OpenStudy (swissgirl):

What are the first 3 terms of the Taylor expansion for the function \(f(x)=e^{2x}\)

OpenStudy (swissgirl):

\(f(x)=e^{2x}\) \(f'(x)=2e^{2x}\) \(f''(x)=4e^{2x}\)

OpenStudy (perl):

the taylor series is , does it say about x = 0 ? i assume it is

OpenStudy (swissgirl):

I am assuming that x=0 too

OpenStudy (swissgirl):

I am not sure why we assume x=0 in this case but whtvr

OpenStudy (perl):

sum f^(n) (a) (x-a)^n / n! = f(a) + f'(a) (x-a) + f '' (a) (x-a)^2/2! + ...

OpenStudy (perl):

we dont have to, it just makes it much cleaner

OpenStudy (perl):

and by cleaner i mean simpler , since x - 0 = x , instead of x - 1 or x - 2 , etc

OpenStudy (perl):

so now lets plug in f(0) + f'(0) *x + f ' ' (0) * x^2/2 + ... , since we want just the first three terms

OpenStudy (swissgirl):

So our answer wld be 1+2+2?

OpenStudy (perl):

e^(2*0) + 2*e^(2*0) * x + 4*e^(2*0) *x^2/2

OpenStudy (perl):

, remember a taylor series is approximating a complicated function (here exponential) with a less complicated function (polynomial)

OpenStudy (perl):

1 , 2x , 4x^2/2

OpenStudy (perl):

ok?

OpenStudy (swissgirl):

Yaaaa I think I get it

OpenStudy (swissgirl):

I get it lol just need to know how to enter it into a program. Just gotta fiddle around. THANKS

OpenStudy (swissgirl):

THANKKSSSSS :))))))))

OpenStudy (swissgirl):

Ok figured it out. Thanks for ur help

OpenStudy (perl):

:D

OpenStudy (perl):

where is my medal?

OpenStudy (perl):

thanks

OpenStudy (swissgirl):

U got ur medal even b4 i got my answer :P